OpenStudy (anonymous):

A man 50 years old has 8 sons born of equal intervals. The sum of the ages of the father and sons is 186. What is the age of the eldest son if the youngest is 3 years old.

OpenStudy (anonymous):

First of all we find the sum of ages of the sons=186-50=136 Then we have an arithmetic progression

OpenStudy (anonymous):

we note with a1=3 the age of the youngest son=>a1=3;r=ratio; 3+a2+a3+a4+a5+a6+a7+a8=136=>3+3+r+3+2r+3+3r+...+3+7r=136=> =>24+28r=136=>28r=112=>r=4; a8=a1+7r=3+7*4=31 =>The age of the eldest son is 31

OpenStudy (anonymous):

Thank you :)